OpenStudy (superman499):

The y-intercept is the value for ______ at the point on the graph where ___=0

OpenStudy (michele_laino):

hint: the y-intercept of this equation: \[\large y = \frac{2}{5}x - 1\] is the value of y,at x=0: \[\large y = \frac{2}{5} \times 0 - 1 = - 1\]

OpenStudy (superman499):

thanks man i appreciate it

OpenStudy (michele_laino):

thanks!

OpenStudy (superman499):

ur welcome